The larger of two numbers is eight more than three times the smaller number. The sum of the two numbers is fifty-two. Find the two numbers.
Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

x and 52-x are the numbers
x=3(52-x)+8
x=156-3x+8
4x=164
x=41
52-x=11

Answer by greenestamps(13198)



In the solution from the other tutor, the problem is set up using the second given statement (the sum of the two numbers is 52) and the problem is then solved using an equation based on the first given statement.

That solution method is fine -- but not the most obvious method. Most students would choose to set up the problem using the first statement and then use the second statement to make an equation to solve to find the answer.

The larger of two numbers is eight more than three times the smaller number:

let the smaller number be x
then the larger number is 3x+8

The sum of the two numbers is fifty-two:

x+3x+8=52
4x+8=52
4x=44
x=11

ANSWER:

The two numbers are
x=11
3x+8=33+8=41

Both methods are equally good. When you have similar problems to solve, try both methods and see what works best for you.